# Sim-to-real Transfer of Deep Reinforcement Learning Agents for Online Coverage Path Planning

### v1

@@ Abstract
Sim-to-real transfer presents a difficult challenge, where models trained in simulation are to be deployed in the real world. The distribution shift between the two settings leads to biased representations of the perceived real-world environment, and thus to suboptimal predictions. In this work, we tackle the challenge of sim-to-real transfer of reinforcement learning (RL) agents for coverage path planning (CPP). In CPP, the task is for a robot to find a path that visits every point of a confined area. Specifically, we consider the case where the environment is unknown, and the agent needs to plan the path online while mapping the environment. We bridge the sim-to-real gap through a semi-virtual environment with a simulated sensor and obstacles, while including real robot kinematics and real-time aspects. We investigate what level of fine-tuning is needed for adapting to a realistic setting, comparing to an agent trained solely in simulation. We find that a high model inference frequency is sufficient for reducing the sim-to-real gap, while fine-tuning degrades performance initially. By training the model in simulation and deploying it at a high inference frequency, we transfer state-of-the-art results from simulation to the real domain, where direct learning would take in the order of weeks with manual interaction, i.e., would be completely infeasible.
1 Introduction
@@
A practical limitation for training machine learning models is the need for large amounts of real-world data, which is tedious and time-consuming to collect. In particular, training reinforcement learning agents for robotics from scratch requires access to the robot for the full duration of the training process. Meanwhile, during the early training phase, the agent is more likely to make mistakes that may damage the hardware or require human supervision for resetting the episodes. Instead, learning in simulation presents an attractive alternative. However, due to differences in the dynamics between the simulation and the real world, transferring a model from simulation to reality is challenging. Prior work reduce the sim-to-real gap by improving the simulation, e.g. by injecting realistic noise [19, 39], applying domain randomization [22, 29], or through meta learning [3, 23]. Our goal is to transfer, for the task of coverage path planning (CPP), the state-of-the-art performance of models trained in simulation to real environments.
@@
In coverage path planning, the task is to find a path that covers all of the free space of an environment. If the environment is known, an optimal path can be planned offline [15]. If it is unknown, the path has to be planned online during mapping of the environment, e.g. by a robot, and an optimal path cannot be found in the general case [9]. CPP has found its uses in many robotic applications, such as lawn mowing [5], vacuum cleaning [36], search-and-rescue [17], and exploration [33].
@@
When it comes to training reinforcement learning models in real time on physical robots, additional considerations need to be accounted for compared to training in simulation. (1) There exists a mismatch between the real and simulated robot kinematics, dynamics, and sensing, such as slippage and noisy localization. This leads to different transition probabilities in the real world, where the predicted actions based on training in simulation may be suboptimal. (2) Due to inertia and latencies in the system, the dynamics are non-Markovian [13]. This violates the common assumption that the environment can be modelled as a Markov decision process satisfying the Markov property. (3) Since the robot keeps moving during the various computations in the training process, the state observed by the agent is not perfectly aligned with the true state. This introduces a delay, where the agent predicts actions based on outdated observations of the environment.
@@
To smoothen the transition into the real world, we use a real robot in a semi-virtual environment, utilizing a highly accurate positioning system in an indoor planar setting, with a simulated sensor and simulated obstacles. This introduces the previously mentioned real-time aspects of reinforcement learning in a controlled manner, while allowing flexibility in creating training environments without having to physically construct them. By fine-tuning a model in the real environment, it approaches a fully realistic setting, although additional real-world aspects such as uneven terrain and localization noise introduced by e.g. a SLAM method are still to be included.
@@
To reduce latency in the training process, we perform model updates in parallel with the action and state selection [38], and perform all computations on on-board hardware. We utilize soft actor-critic learning [12] for its sample efficiency and its efficacy on continuous action spaces. To account for the non-Markovian dynamics in a lightweight manner, we include past actions in the observation space. Finally, to reduce the mismatch between the simulated and real kinematics, we measure the real-world linear and angular accelerations as well as action delay, and incorporate them into the simulation. We find that fine-tuning the model initially degrades performance, and a long training time is likely required to adapt to the distribution shift. Meanwhile, deploying the model at a high inference frequency without fine-tuning proved to be sufficient for sim-to-real transfer.
@@
  • We show that this division enables a transfer of the state-of-the-art RL policy for CPP from simulation to the real robot.
@@
  • We evaluate the effects of the time step size and the fine-tuning on the performance of the real robot.
2 Related Work
@@
This paper relates to coverage path planning, transferring models from simulation to the real world, and training reinforcement learning online in real time. We summarize the related work below.
@@
Coverage path planning methods can roughly be categorized as planning-based or learning-based. Planning-based methods include decomposition methods, which divide the area into cells based on e.g. boustrophedon cellular decomposition (BCD) [7] or Morse decomposition [1], where each cell is covered with a pre-defined pattern. Grid-based methods, such as Spiral-STC [8] and the backtracking spiral algorithm (BSA) [10], discretize the area into even smaller cells, e.g. squares, with a similar size to the agent. Then, a path is planned based on motion primitives connecting adjacent cells, i.e. to move up, down, left, or right in the grid, such that each cell is visited at least once. Frontier-based methods, on the other hand, plan a path to a chosen point on the frontier, i.e. on the boundary between covered and non-covered regions. The choice of frontier point can be based on different aspects, such as the distance to the agent [34], the path of a rapidly exploring random tree (RRT) [30] or the gradient in a potential field [37]. Learning-based methods apply machine learning techniques, typically in combination with planning-based methods, to find coverage paths. Reinforcement learning is the most popular machine learning paradigm, due to the sequential nature of the task. Chen et al. [6] use RL to find the order in which to cover the cells generated by BCD. Discrete methods [25, 20] use RL to learn which motion primitives to perform. RL has also been combined with frontier-based methods, either for predicting the cost of each frontier point [24], or for learning the control signals to navigate to a chosen point [14]. In contrast, Jonnarth et al. [18] propose to learn continuous control signals end-to-end from sensor data and a map of the environment. In our work, we adopt this approach to transfer an agent trained in simulation to a real setting.
@@
Sim-to-real transfer. Transferring from simulation to the real world is challenging due to mismatch in both sensing and actuation [40]. Prior work has approached this challenge by different means. Domain randomization has been utilized to randomize physical parameters in simulation, such as mass and joint damping [22], or to randomize textures and lightning in the image domain [29]. Closely related, other works introduce perturbations in the sensing and actuation [39]. Meta learning methods aim to quickly adapt to new unseen tasks from a wide variety of training task, such as adapting to the real world from simulation [3, 23]. Another approach is to learn from expert demonstrations through imitation learning [35], which has previously been applied to coverage path planning [14]. When it comes to robot control, Niroui et al. [24] deploy an RL policy for CPP trained in simulation on a differential drive robot without fine-tuning. The policy predicts the next frontier node, where a separate non-learned module navigates to it, thus being less affected by misaligned kinematics between simulation and reality. Kaufmann et al. [19] transfer a lightweight RL control policy using a pre-trained perception system for first-person-view drone racing, utilizing empirical noise models to improve the fidelity of the simulator. They collect perception and dynamics residuals in the real world based on a highly accurate positioning system, which they use to augment the simulation where the control policy is fine-tuned. In contrast to these works, we fine-tune our model online in the real world.
@@
Online RL. Compared to turn-based and simulated environments, where the time for action selection and policy updates are assumed to be zero and the next state can be advanced to in an instant, online RL in the real world presents new challenges where these assumptions do not hold. The core idea in the literature for accounting for this discrepancy is to parallelize aspects of the environment interactions and the model training. Bakker et al. [4] explore quasi-online reinforcement learning, where a model of the environment is built online and in parallel to training the policy based on the built environment model. Ramstedt and Pal [28] propose real-time reinforcement learning, where the action selection occurs in parallel with the state selection. Concretely, given the current state and action to be taken, the agent and environment compute the next set of state and action concurrently. This approach takes into account the time for action selection by delaying the observed state by one time step. However, it does not consider the time for model updates, which is typically much larger. Other works [11, 31] distribute the model training for soft actor-critic (SAC) [12] to a separate system, while performing only the lightweight action and state selection computations on the on-board target hardware. This allows the data collection to be executed independently from the model updates, where the updated policy and latest data are periodically synchronized between the two systems. Yuan and Mahmood [38] employ a similar approach, but they run everything on a single edge device. They divide the training process into separate processes for the action selection, state selection, batch sampling, and gradient computation. This ensures that the data collection is unencumbered by the policy updates, while new training batches are sampled in parallel to allow gradient computations to run without interruption. We follow this approach. However, in our experimental setup, the action selection and batch sampling were much faster than the state selection and gradient computations, so we only use two threads in order to reduce complexity and communication overhead.
3 Background
@@
In this section we first formulate the CPP problem as a Markov decision process, and then briefly describe the approach by Jonnarth et al. [18], which we use to train RL agents for CPP.
3.1 Problem Formulation
@@
We formulate the task as a partially observable Markov decision process (POMDP), consisting of an agent that predicts actions in an environment such as to maximize a reward function. In discrete time steps $t\in[1..T]$, the agent predicts actions $a_{t}\sim\pi(a_{t}|o_{t})$ according to policy $\pi$, based on observations $o_{t}$ of the state $s_{t}$. The policy is a neural network that predicts continuous control signals $a_{t}$ for the agent. Subsequently, the agent receives a new observation and reward $o_{t},r_{t}\sim p(o_{t},r_{t}|s_{t},a_{t-1})$ from the environment, where $s_{t}\sim p(s_{t}|s_{t-1},a_{t-1})$. The goal for the agent is to maximize the expected discounted reward $\mathbb{E}(\sum_{t=1}^{T}\gamma^{t}r_{t})$ with discount factor $\gamma$.
In our particular problem, the state $s_{t}$ includes the full environment geometry with all obstacles, the set of points that have been covered, as well as the pose of the agent. However, the agent only has information about what it has observed, so the observation $o_{t}$ consists of the part of the environment that has been mapped until time step $t$, along with covered points, agent pose, and sensor data. In the setting with a real robot, the first order Markovian property is not fulfilled, due to the dynamics including inertia and momentum. The usual approach to augment previous states in the current one becomes infeasible if the motion is implicitly represented by the agent-centered environment geometry. Thus higher-order effects lead to model errors that need correction in subsequent steps.
3.2 Learning Coverage Paths with Reinforcement Learning
Following Jonnarth et al. [18], we train a scale-grouped CNN (SGCNN) to predict control signals for CPP based on coverage, obstacles, and frontier maps in multiple scales, and apply dense coverage and total variation rewards. We briefly describe these components in this section.
Map observation. To represent the environment, we keep track of the covered region and detected obstacles in the form of global grid-like 2D maps. In order to represent large regions as input to the agent, we feed it multiple local maps with varying scales, each with the same grid size, where the maps contain increasingly larger crops of the area with decreasing resolution. The finest scale shows detailed information near the agent, while the coarsest scale contains the largest area at a low resolution. In addition to coverage and obstacle maps, we also encode a multi-scale frontier map as in [18]. The frontier is the boundary between covered and non-covered space, which allows the agent to easily plan paths towards non-covered regions. All maps are represented as egocentric maps, i.e. they are aligned such that the agent is in the center facing upwards.
@@ 3.2 Learning Coverage Paths with Reinforcement Learning
Action space. We consider a differential drive wheeled robot, which is controlled by two separately driven wheels. The agent predicts the linear and angular velocities $v$ and $\omega$ for the robot, which are converted to angular velocities $\omega_{R}$ and $\omega_{L}$ for the right and left wheels. These are given by
@@
Reward function. In addition to a coverage reward term, Jonnarth et al. [18] found it necessary to include a total variation (TV) reward that penalizes variations in the coverage map, such as excessive boundaries. By reducing the total variation, the agent learns to cover the area in a more compact manner, reducing leftover holes in the coverage map, and thus, allows for complete coverage. In total, we use four reward terms; namely, a coverage reward $R_{\mathrm{area}}$, a TV reward $R_{\mathrm{TV}}$, a collision reward $R_{\mathrm{coll}}$, and a constant reward $R_{\mathrm{const}}$. The coverage reward is written as
@@
Finally, the agent is penalized by a negative reward $R_{\mathrm{coll}}$ when it collides with an obstacle, and a negative constant reward $R_{\mathrm{const}}$ is given each time step.
Model architecture. To process the multi-scale maps together with the sensor data, we use a scale-grouped convolutional neural network (SGCNN) [18]. It consists of a map feature extractor $g_{m}$, a sensor feature extractor $g_{s}$, and a fusing module $g_{f}$. The map feature extractor groups the maps by scale and processes them separately using grouped convolutions, while the sensor feature extractor and fusing module consist of fully connected layers. The control signal predictions are computed as
@@
where $M_{c}$, $M_{o}$, and $M_{f}$ are the multi-scale coverage, obstacle, and frontier maps respectively, and $S$ is a vector of the observed lidar distance measurements.
4 Transferring CPP Agents from Simulation to the Real World
The approach described in the previous section produces CPP-agents with state-of-the-art performance in simulation. Obviously, one could train the same agent directly in the real environment, which would require training with manual interaction in the order of weeks. In this section, we describe our proposed approach, which allows to transfer the model trained during simulation to the real robot.
4.1 Learning CPP on the Real System
Even if we will not train the agent from scratch in the real setting, we need to be able to fine-tune the transferred system by RL. In common RL-libraries, such as stable-baselines3 [27], tianshou [32], and spinningup [2], the data collection and model updates are performed serially. While this is feasible in simulation and turn-based environments where the environment can be paused during model updates and time can be advanced instantly to the next state after action selection, it is not practical for real-time robotic applications where the agent is trained online.
In this case, we need to wait after action selection to observe the next state, while the execution remains idle. After the environment interaction, a batch is sampled from the replay buffer, gradients are computed, and the model is updated. However, this takes time, which is not negligible, especially on low-performance edge devices and embedded systems. During the gradient update step, the robot keeps following its previous action. As a result, the agent ends up in a state which deviates from the recorded one used for action selection. Following previous work [11, 38], we perform the model updates in parallel with the data collection, utilizing computational resources which would otherwise remain idle while waiting for the next state.
@@ 4.1 Learning CPP on the Real System
Table 1: Measured times of the RL-step on the real system. Timing for (2) is an application dependent design choice. The overhead effectively becomes part of (1) resulting in an action delay of 50 ms.
@@
Since the action selection and batch sampling were fast compared to the state selection and gradient computations, we only use two threads to reduce overhead and complexity. Our computational graph is shown in Figure 1, which consists of an environment interaction thread (left) and a model update thread (right). Since both threads interact with the replay buffer, we use a mutex to avoid conflicts. Luckily, the the environment interaction thread adds an entry at the end of the cycle, while the model update thread samples from it early, so the there is a low chance that they block each other.
Figure 1: During online training we perform environment interactions and model training in parallel.
Since the model update thread modifies the weight tensors during a large portion of the runtime, i.e. both during the backward pass to update the gradients and during the optimizer step to update the parameters, a mutex is not feasible as it would block execution too often. Instead, we keep a copy of the model, which is only accessed by the model update thread. We then synchronize the weights when both threads have finished.
With this approach, both the environment interaction and the model update can be performed multiple times before model synchronization. This is useful when the computation time for the model update exceeds that of the action and state selection. In this case, the number of environment interaction steps should be chosen to match the computation time for the model update. Meanwhile, if the model update is fast, multiple updates can be performed during one environment interaction step.
Apart from the four main computational steps, the time for simulating the sensor, creating the observation, and synchronizing the weights results in additional overhead. This effectively becomes part of the action selection step as the overhead delays the observed state, and results in an action delay.
4.2 Smoothening the Sim-to-Real Gap
With the approach in the previous section, we could directly move to the real setting. However, this would make training and fine-tuning cumbersome as manual interaction would be required. In order to fully automatically (continue to) train the system, we propose to use a semi-virtual setup, see Figure 2.
@@ 4.2 Smoothening the Sim-to-Real Gap
In this setup we use the real robot with its kinematics and dynamics, but simulate the sensor on the robot, a lidar sensor, and localize the robot using a positioning system. The environment, including obstacles and the result from, e.g., mowing, is visualized by a projection onto the ground.
As mentioned above, this semi-virtual setup enables fully automatic RL with the real robot. In particular if the robot would drive towards one of the walls, both the robot and the environment can be moved back to the middle of the room. Furthermore, the setup can also be used for fully automatic benchmarking of the learned model.

### v2

@@ Abstract
Sim-to-real transfer presents a difficult challenge, where models trained in simulation are to be deployed in the real world. The distribution shift between the two settings leads to biased representations of the dynamics, and thus to suboptimal predictions in the real-world environment. In this work, we tackle the challenge of sim-to-real transfer of reinforcement learning (RL) agents for coverage path planning (CPP). In CPP, the task is for a robot to find a path that covers every point of a confined area. Specifically, we consider the case where the environment is unknown, and the agent needs to plan the path online while mapping the environment. We bridge the sim-to-real gap through a semi-virtual environment, including a real robot and real-time aspects, while utilizing a simulated sensor and obstacles to enable environment randomization and automated episode resetting. We investigate what level of fine-tuning is needed for adapting to a realistic setting, comparing to an agent trained solely in simulation. We find that a high inference frequency allows first-order Markovian policies to transfer directly from simulation, while higher-order policies can be fine-tuned to further reduce the sim-to-real gap. Moreover, they can operate at a lower frequency, thus reducing computational requirements. In both cases, our approaches transfer state-of-the-art results from simulation to the real domain, where direct learning would take in the order of weeks with manual interaction, that is, it would be completely infeasible.
Introduction
@@
A practical limitation for training machine learning models is the need for large amounts of real-world data, which is tedious and time-consuming to collect. In particular, training reinforcement learning (RL) agents for robotics from scratch requires access to the robot for the full duration of the training process. Meanwhile, during the early training phase, the agent is more likely to make mistakes that may damage the hardware or require human supervision for resetting the episodes. Instead, learning in simulation presents an attractive alternative (Jonnarth, Zhao, and Felsberg 2024). However, due to differences in the dynamics between the simulation and the real world, transferring a model from simulation to reality is challenging. Prior work reduce the sim-to-real gap by improving the simulation, e.g. by injecting realistic noise (Kaufmann et al. 2023; Zhao et al. 2020), applying domain randomization (Muratore et al. 2021; Tobin et al. 2017), or through meta learning (Arndt et al. 2020; Nagabandi et al. 2019). Our goal is to transfer, for the task of coverage path planning (CPP), the state-of-the-art performance of models trained in simulation to real environments, by fine-tuning in a realistic setting.
@@
In CPP, the task is to find a path that covers all of the free space of an environment. If the environment is known, an optimal path can be planned offline (Huang 2001). If it is unknown, the path has to be planned online during mapping of the environment, e.g. by a robot, and an optimal path cannot be found in the general case (Galceran and Carreras 2013). CPP has found its uses in many robotic applications, such as lawn mowing (Cao, Huang, and Hall 1988), vacuum cleaning (Yasutomi, Yamada, and Tsukamoto 1988), search-and-rescue (Jia et al. 2016), and exploration (Xu et al. 2022).
@@
When training reinforcement learning models in real time on physical robots, additional considerations need to be accounted for compared to training in simulation. (1) There exists a mismatch between the real and simulated robot kinematics, dynamics, and sensing, such as slippage and noisy localization. This leads to different transition probabilities in the real world, where the predicted actions based on simulation may be suboptimal. (2) Due to inertia and latencies in the system, the dynamics are non-Markovian (Haarnoja et al. 2018b). This violates the common assumption that the Markov property holds. (3) Since the robot keeps moving during the various computations in the training process, the state observed by the agent is not perfectly aligned with the true state. This introduces a delay, where the agent predicts actions based on outdated observations of the environment. In this work, we address these problems.
@@
To smoothen the transition into the real world, we use a real robot in a semi-virtual environment, utilizing a highly accurate positioning system in an indoor planar setting, with a simulated sensor and obstacles. This introduces the previously mentioned real-time aspects of reinforcement learning in a controlled manner, while allowing flexibility in utilizing automated episode resetting and varying training environments without having to physically construct them. By fine-tuning and evaluating CPP policies in this semi-virtual setting, we can draw conclusions about how we expect similar policies to generalize to fully realistic environments.
@@
To reduce latency in the training process, we perform model updates in parallel with the action and state selection (Yuan and Mahmood 2022), and perform all computations on on-board hardware. We utilize soft actor-critic learning (Haarnoja et al. 2018a) for its sample efficiency and its efficacy on continuous action spaces. To account for the non-Markovian dynamics in a lightweight manner, we include past actions in the observation space. Finally, to reduce the mismatch between the simulated and real kinematics, we measure the real-world linear and angular accelerations as well as action delay, and incorporate them into the simulation. We find that a high inference frequency enables first-order Markovian policies to transfer to a real setting. By introducing delays and action observations, higher order Markovian models can be fine-tuned to further reduce the sim-to-real gap. Moreover, these models can operate at a lower frequency, thus reducing the computational requirements for deployed systems.
@@
  • We show that this division enables a transfer of state-of-the-art RL policies for CPP from simulation to the real robot.
@@
  • We evaluate the effects of inference frequency and fine-tuning on the performance of the real robot.
Related Work
@@
This paper relates to coverage path planning, transferring models from simulation to the real world, and online RL in real time. We summarize the related work below.
@@
Coverage path planning methods can roughly be categorized as planning-based or learning-based. Planning-based methods include decomposition methods, which divide the area into cells based on e.g. boustrophedon cellular decomposition (BCD) (Choset and Pignon 1998) or Morse decomposition (Acar et al. 2002), where each cell is covered with a pre-defined pattern. Grid-based methods, such as Spiral-STC (Gabriely and Rimon 2002) and the backtracking spiral algorithm (BSA) (Gonzalez et al. 2005), discretize the area into even smaller cells, e.g. squares, with a similar size to the agent. Then, a path is planned based on motion primitives connecting adjacent cells, i.e. to move up, down, left, or right on the grid, such that each cell is visited at least once. Frontier-based methods plan a path to a chosen point on the frontier, i.e. on the boundary between covered and non-covered regions. The choice of frontier point can be based on different aspects, such as the distance to the agent (Yamauchi 1997), the path of a rapidly exploring random tree (RRT) (Umari and Mukhopadhyay 2017) or the gradient in a potential field (Yu et al. 2021). Learning-based methods apply machine learning techniques, typically in combination with planning-based methods, to find coverage paths. Reinforcement learning is the most popular paradigm due to the sequential nature of the task. Chen et al. (2019) use RL to find the order in which to cover the cells generated by BCD. Discrete methods (Piardi et al. 2019; Kyaw et al. 2020) use RL to learn which motion primitives to perform. RL has also been combined with frontier-based methods, either for predicting the cost of each frontier point (Niroui et al. 2019), or for learning the control signals to navigate to a chosen point (Hu et al. 2020). In contrast, Jonnarth, Zhao, and Felsberg (2024) propose to learn continuous control signals end-to-end from a built map and sensor data. In our work, we adopt this approach to transfer agents from simulation to a real setting. To the best of our knowledge, we are the first to do so for end-to-end RL policies in CPP.
@@
Sim-to-real transfer. Transferring from simulation to the real world is challenging due to mismatch in both sensing and actuation (Zhao, Queralta, and Westerlund 2020). Prior work has approached this challenge by different means. Domain randomization has been utilized to randomize physical parameters in simulation, such as mass and joint damping (Muratore et al. 2021), or to randomize textures and lighting in the image domain (Tobin et al. 2017). Other works introduce perturbations in the sensing and actuation (Zhao et al. 2020). Meta learning methods aim to quickly adapt to new unseen tasks from a wide variety of training task, such as adapting to the real world from simulation (Arndt et al. 2020; Nagabandi et al. 2019). Another approach is to learn from expert demonstrations through imitation learning (Yan et al. 2017), which has previously been applied to coverage path planning (Hu et al. 2020). When it comes to robot control, Niroui et al. (2019) deploy an RL policy for CPP trained in simulation on a differential drive robot without fine-tuning. The policy predicts the next frontier node, where a separate non-learned module navigates to it, thus being less affected by misaligned kinematics between simulation and reality. Kaufmann et al. (2023) transfer a lightweight RL control policy using a pre-trained perception system for first-person-view drone racing, utilizing empirical noise models to improve the fidelity of the simulator. They collect perception and dynamics residuals in the real world based on a highly accurate positioning system, which they use to augment the simulation. In contrast to these works, we fine-tune our model online in the real world.
@@
Online RL. Compared to turn-based and simulated environments, where the time for action selection and policy updates are assumed to be zero and the next state can be advanced to in an instant, online RL in the real world presents new challenges where these assumptions do not hold. The core idea in the literature for accounting for this discrepancy is to parallelize aspects of the environment interactions and the model training. Bakker et al. (2006) explore quasi-online reinforcement learning, where a model of the environment is built online and in parallel to training the policy based on the built environment model. Ramstedt and Pal (2019) propose real-time reinforcement learning, where the action selection occurs in parallel with the state selection. Concretely, given the current state and action to be taken, the agent and environment compute the next set of state and action concurrently. This approach takes into account the time for action selection by delaying the observed state by one time step. However, it does not consider the time for model updates, which is typically much larger. Other works (Haarnoja et al. 2019; Wang, Vasan, and Mahmood 2023) distribute the model training for soft actor-critic (SAC) (Haarnoja et al. 2018a) to a separate system, while performing only the lightweight action and state selections on the on-board target hardware. This allows the data collection to be executed independently from the model updates, where the policy and replay buffer are periodically synchronized. Yuan and Mahmood (2022) employ a similar approach, but they run everything on a single edge device. They divide the training process into separate processes for the action selection, state selection, batch sampling, and gradient computation. We follow this approach. However, in our experimental setup, the action selection and batch sampling were much faster than the state selection and gradient computations, so we only use two threads in order to reduce complexity and communication overhead.
Background
@@
In this section, we first formulate the CPP problem as a partially observable Markov decision process, and then briefly describe the approach by Jonnarth, Zhao, and Felsberg (2024), which we use to train RL agents for CPP.
Problem Formulation
@@
We formulate the task as a partially observable Markov decision process (POMDP), consisting of an agent that predicts actions in an environment such as to maximize a reward function. In discrete time steps $t\in\{1,...,T\}$, the agent predicts actions $a_{t}\sim\pi(a_{t}|o_{t})$ according to policy $\pi$, based on observations $o_{t}$ of the state $s_{t}$. The policy is a neural network that predicts continuous control signals $a_{t}$ for the agent. Subsequently, the agent receives a new observation and reward $o_{t},r_{t}\sim p(o_{t},r_{t}|s_{t},a_{t-1})$ from the environment, where $s_{t}\sim p(s_{t}|s_{t-1},a_{t-1})$. The goal for the agent is to maximize the expected discounted reward $\mathbb{E}(\sum_{t=1}^{T}\gamma^{t}r_{t})$, with discount factor $\gamma$.
In our particular problem, the state $s_{t}$ includes the full environment geometry with all obstacles, the set of points that have been covered, as well as the pose of the agent. However, the agent only has information about what it has observed, so the observation $o_{t}$ consists of the part of the environment that has been mapped until time step $t$, along with covered points, agent pose, and sensor data. In the setting with a real robot, the first order Markovian property is not fulfilled, due to the dynamics including inertia and momentum. The usual approach to augment previous states in the current one becomes infeasible if the motion is implicitly represented by the agent-centered environment maps. Thus higher-order effects lead to model errors that need correction in subsequent steps.
Reinforcement Learning of Coverage Paths
In this section, we summarize the approach presented by Jonnarth, Zhao, and Felsberg (2024) for learning CPP with RL, in terms of observation space, action space, reward function, and model architecture.
Map observation. In order to represent the environment, covered regions and detected obstacles are recorded in the form of global grid-like 2D maps. To represent large regions as input to the agent, multiple local maps are used with varying scales, each with the same grid size. The maps contain increasingly larger crops of the local area, with decreasing resolution. The finest scale shows detailed information near the agent, while the coarsest scale contains the largest area at a low resolution. In addition to coverage and obstacle maps, multi-scale frontier maps are also recorded. The frontier is the boundary between covered and non-covered space, which allows the agent to easily plan paths towards non-covered regions. All maps are represented as egocentric maps, that is, they are aligned such that the agent is in the center facing upwards.
@@
Action space. We consider a differential drive wheeled robot, which is controlled by two separately driven wheels. The agent predicts the linear and angular velocities $v$ and $\omega$ for the robot. In contrast to the cited work, we convert them to angular velocities $\omega_{R}$ and $\omega_{L}$ for the right and left wheels,
@@
Reward function. In addition to a coverage reward term, Jonnarth, Zhao, and Felsberg (2024) found it necessary to include a total variation (TV) reward that penalizes variations in the coverage map, such as excessive boundaries. By reducing the total variation, the agent learns to cover the area in a more compact manner, reducing leftover holes in the coverage map, and thus, allows for complete coverage. In total, four reward terms are used: a coverage reward $R_{\mathrm{area}}$, a TV reward $R_{\mathrm{TV}}$, a collision reward $R_{\mathrm{coll}}$, and a constant reward $R_{\mathrm{const}}$. The coverage reward is written as
@@
Finally, the agent is penalized by a negative obstacle collision reward $R_{\mathrm{coll}}$ on impact, and a constant negative reward $R_{\mathrm{const}}$ is given each time step. Each episode is terminated when a goal coverage has been reached, or after $\tau=1000$ consecutive steps without new ground being covered.
Model architecture. To process the multi-scale maps together with the sensor data, a scale-grouped convolutional neural network (SGCNN) is used, which has shown to be superior to a standard CNN or MLP (Jonnarth, Zhao, and Felsberg 2024). It consists of a map feature extractor $g_{m}$, a sensor feature extractor $g_{s}$, and a fusing module $g_{f}$. The map feature extractor groups the maps by scale and processes them separately using grouped convolutions, while the sensor feature extractor and fusing module consist of fully connected layers. The control signal predictions are given by
@@
where $M_{c}$, $M_{o}$, and $M_{f}$ are the multi-scale coverage, obstacle, and frontier maps respectively, and $S$ is a vector of the observed normalized lidar distance measurements.
Transferring CPP Agents to the Real World
The approach described in the previous section produces CPP agents with state-of-the-art performance in simulation. Obviously, one could train the same agent directly in the real environment. However, this would require training with manual interaction in the order of weeks to months. In this section, we describe our proposed approach, which allows to transfer the model trained in simulation to the real robot.
Learning CPP on the Real System
Even if we will not train the agent from scratch in the real setting, we need to be able to fine-tune the transferred system by RL. In common RL-libraries, such as stable-baselines3 (Raffin et al. 2021), tianshou (Weng et al. 2022), and spinningup (Achiam 2018), the data collection and model updates are performed serially. While this is feasible in simulation and turn-based environments where the environment can be paused during model updates and time can be advanced instantly to the next state after action selection, it is not practical for real-time robotic applications where the agent is trained online.
In this case, we need to wait after action selection to observe the next state, while the execution remains idle. After the environment interaction, a batch is sampled from the replay buffer, gradients are computed, and the model is updated. However, this takes time, which is not negligible, especially on low-performance edge devices and embedded systems. During the gradient update step, the robot keeps following its previous action. As a result, the agent ends up in a state which deviates from the recorded one used for action selection. Following previous work (Haarnoja et al. 2019; Yuan and Mahmood 2022), we perform the model updates in parallel with the data collection, utilizing computational resources which would otherwise remain idle while waiting for the next state.
Figure 1: During online training we perform environment interactions (top) and model training (bottom) in parallel.
@@
Table 1: Measured times of the RL-step on the real system. Timing for (2) is application dependent, and chosen based on the optimal training step size. The overhead effectively becomes part of (1) resulting in an action delay of 50 ms.
Since the action selection and batch sampling were fast compared to the state selection and gradient computations, we only use two threads to reduce overhead and complexity. Our computational graph is shown in Figure 1, which consists of an environment interaction thread (top) and a model update thread (bottom). Since both threads interact with the replay buffer, we use a mutex to avoid conflicts. Luckily, the the environment interaction thread adds an entry at the end of the cycle, while the model update thread samples from it early, so the there is a low chance that they block each other.
Since the model update thread modifies the weight tensors during a large portion of the runtime, i.e. both during the backward pass to update the gradients and during the optimizer step to update the parameters, a mutex is not feasible as it would block execution too often. Instead, we keep a copy of the model, which is only accessed by the model update thread. The weights are synchronized when both threads have finished.
With this approach, both the environment interaction and the model update can be performed multiple times before model synchronization. This is useful when the computation time for the model update exceeds that of the action and state selection. In this case, the number of environment interaction steps should be chosen to match the computation time for the model update. Meanwhile, if the model update is fast compared to the environment interaction, multiple updates can be performed during one environment interaction step, which was the case in our experiments.
Apart from the four main computational steps, the time for simulating the sensor, creating the observation, and synchronizing the weights results in additional overhead. This effectively becomes part of the action selection as the overhead delays the observed state, and results in an action delay.
Smoothening the Sim-to-Real Gap
With the approach in the previous section, we could directly move to a real setting in-the-wild. However, this would make training and fine-tuning cumbersome as manual interaction would be required to supervise the training process. In order to fully automatically (continue to) train the system, we propose to use a semi-virtual setup, see Figure 2. In this setup, we use the real robot with its kinematics and dynamics, but simulate its lidar sensor and obstacles, and localize the robot using a positioning system. The environment, including obstacles and coverage, e.g. from mowing, is visualized by a projection onto the ground.
The main purpose of using such an environment is to evaluate, in a controlled manner, the ability of the learning algorithm to generalize to a real robot in a real-time setting. Furthermore, the considered RL approach is robust to sensing and localization noise, and can be extended to the case of unknown pose, e.g. by estimating the pose using an off-the-shelf SLAM method (Jonnarth, Zhao, and Felsberg 2024). Thus, using our semi-virtual environment, we can draw conclusions about how we expect similar policies to generalize to fully realistic settings.
@@
As mentioned, this semi-virtual setup enables fully automatic RL with the real robot. In particular, if the robot would drive towards one of the walls, both the robot and the environment can be moved back to the middle of the room. Furthermore, since we can automatically change or generate new environments each episode, we can train general CPP policies for completely unseen environments, and not just for a single fixed environment. Finally, the setup can also be used for fully automatic benchmarking of the learned model.